### v1

# Improving Zero-noise Extrapolation for Quantum-gate Error Mitigation using a Noise-aware Folding Method

Abstract
The current thousand-qubit processors mark a substantial advance in hardware. Yet, hardware limitations prevent quantum error correction (QEC), necessitating reliance on quantum error mitigation (QEM). Our paper presents a noise-aware folding method that improves Zero-Noise Extrapolation (ZNE) by estimating noiseless values from noisy results. Unlike traditional ZNE methods, which assume a uniform error distribution, our method redistributes the noise using calibration data based on hardware noise models. By employing noise-adaptive compilation and optimizing the qubit mappings, our approach enhances the ZNE accuracy of various quantum computing models. Recalibrating the noise amplification to address the inherent error variations, promises higher precision and reliability in quantum computations. This paper highlights the uniqueness of our method, summarizes noise accumulation, presents the scaling algorithm, and compares the reliability of our method with those of existing models using linear fit extrapolation. Relative to the existing folding methods, our method achieved a 35% improvement on quantum computer simulators and a 26% improvement on real quantum computers compared to existing folding methods, demonstrating the effectiveness of our proposed approach.
preprint: APS/123-QED
I Introduction
Recently, quantum computing (QC) technology has entered the era of noisy-intermediate scale quantum computing (NISQ), highlighted by IBM’s recent release of processors housing over a thousand qubits [1, 2]; meanwhile, robust crosstalk chips using tunable-coupler technology, can operate up to 133 qubits [3, 4, 5]. Despite these advancements, the capacity of current hardware is inadequate for implementing quantum error correction codes but is projected to become adequate by the 2030s [6, 7]. In the meantime, QC systems employ QEM methods, often requiring additional quantum and classical resources as a trade-off to enhance QC output fidelity.
Numerous QEM techniques for alleviating diverse sources of error in QC have been proposed [8]: probabilistic error cancelation for mitigating decoherence  [9, 10], ZNE for imperfect gates, techniques for mitigating measurement errors  [11, 12, 13], dynamical decoupling [14, 15], quantum optimal control [16, 17], randomized compiling [18, 19], Pauli-frame randomization [20, 21], and other techniques [22, 23, 24, 25, 21, 26]. The present study focuses on the well-established ZNE technique.
ZNE was concurrently introduced and extensively demonstrated in various applications involving systems of up to 127 qubits [27, 28]. In ZNE, a quantum program is amplified to different noise levels through gate or pulse-level stretching, which intentionally increases the program’s noise. Subsequently, the amplified results are extrapolated to estimate the noiseless values. Formally, the program is amplified by multiple scale factors $\lambda$. When $\lambda=1$, the program operates at its original error rate. When $\lambda$ $>0$, an additional error, such as an identity unitary combination of the original gate, is introduced. The results obtained at different noise levels of $\lambda$ are collected and extrapolated to $\lambda$ $=0$, effectively eliminating the noise. Therefore, $\lambda$ represents the noise level affecting any physical quantity during quantum computation.
Researchers have proposed various methods for amplifying noise in quantum programs. For instance, the approach described in [29] stretches gate durations to the desired levels using pulse-level control. Under ideal conditions, this stretching does not alter the quantum system’s state, but under noisy conditions, these modifications are classified as errors. Moreover, as this method requires a high degree of abstract control and calibration at the pulse level in quantum computers, it is not easily implementable across most of the existing QC systems. In constrast, unitary folding [30] amplifies the noise using gate-level control, which is available in all gate-based computation model QCs. This technique builds upon a simple concept: replacing the unitary operation $U\rightarrow U\left(U^{\dagger}U\right)$, where $\left(U^{\dagger}U\right)$ is an identity. In an ideal scenario, this replacement remains a $U$ operation, but errors corresponding to U in the quantum system amplify the errors in the $U$ operation within the modified circuit. Giurgica-Tiron et al.  [30] proposed unitary folding with a fold from the left (which folds each unitary gate independently) and at random folding methods (which randomly selects a subset of individual gates randomly as a block of unitary gates and replicates the entire block with $U\left(U^{\dagger}U\right)$. This method requires no knowledge of the QC underlying noise model. Although these scaling methods effectively amplify noise in the original input quantum program, they neglect the sources of error imbalance within the QC system, potentially leading to biased results and errors that may compromise the accuracy of the extrapolation results.
To resolve this problem, we introduce a noise-aware folding method that adjusts the quantum circuit, aiming to balance the error rate across all logical qubits by leveraging the calibration data noise model, which is freely available on select QC systems. This approach comprehensively analyzes the calibration data noise model, thereby redistributing the noise amplification process across the quantum circuit. Through an algorithmic design, our approach strategically adjusts the scaling factor for each gate operation, balancing the distribution of gate error rates across all logical qubits. Addressing the error source imbalances, mitigates bias-induced errors, thereby enhancing the accuracy of ZNE extrapolated results from the quantum program.
Moreover, the approach is adaptable and therefore versatile across various quantum computing models. Seamlessly integrated with gate-based computation models, it comprehensively caters to the inherent complexities of contemporary quantum hardware. The integration of a noise-aware folding method enhances the precision and reliability of quantum computations. The amplification process can be recalibrated to accommodate the intricate error landscapes within quantum systems, promising more accurate and dependable quantum programming, and bridging the gap between theoretical expectations and practical outcomes. The contributions of our study are highlighted below:
  • Our noise-aware folding method is tailored to redistribute noise across quantum circuits, thereby enhancing the accuracy of ZNE by addressing the inherent error variations within quantum systems.
  • We seamlessly integrate our approach with gate-based computation models and optimize the qubit mappings using hardware noise models, providing versatility mitigating bias-induced errors, and enhancing the reliability of quantum computations.
  • Our novel algorithm dynamically adjusts the error rates within the quantum circuit, leveraging calibration data and noise-adaptive compilation methods to ensure a balanced and controlled scaling process.
  • We conduct rigorous experiments across various full noise models and real quantum computers, showcasing the method’s consistent performance in simulations, highlighting the challenges of scaling to larger circuits, and emphasizing the disparity between simulated and real quantum computer executions.
Figure 1: Controlled-not (CNOT) operation as identity insertion of unitary U1 and U2 operations using two qubits.
The remainder of this is organized as follows. Section II.1 briefly discusses the background and related QEM methods using ZNE techniques, which motivated our research. Section III explains the noise-aware folding method, along with its compilation and execution scheme. Section IV analyzes the performance results of our approach and previous methods. Our noise-aware folding method is further discussed in Section V and conclusions are presented in Section VI.
II Background and Related Works
This section briefly provides the fundamentals of our proposed method. We discuss the quantum error mitigation (QEM), noise scaling methods, and noise extrapolation models.
II.1 Quantum Error Mitigation
Quantum computers are susceptible to various sources of noise and imperfections including crosstalk, which can affect the accuracy, reliability, and fidelity of QC results. Utilizing a set of techniques and strategies to minimize quantum computational errors, QEM estimates the true expectation value $\langle A\rangle$ from the approximated value $A$ impacted by errors $\psi$.
$\langle A\rangle=\operatorname{Tr}(A\psi)$ (1)
QEM strategically minimizes the impact of errors at various stages of quantum processes such as ZNE. Ultimately, it aims to optimize the performance of quantum algorithms and facilitate their seamless integration into practical QC environments. As one of the best-estimated QC techniques, ZNE is used for understanding the behavior of quantum algorithms in the absence of noise. In ZNE, quantum computations are run at different noise levels and the outcomes are extrapolated to a hypothetical noiseless scenario (the zero-noise limit). By analyzing the results obtained at various noise levels, we can predict the performance of a quantum algorithm without noise interference. This extrapolation provides insights into the intrinsic properties and capabilities of the quantum algorithm, clarifying its idealized behavior. Essentially, ZNE can separate the impact of noise from the true potential of a quantum algorithm.
II.2 Noise Scaling Methods
ZNE methods measure a given observable at varying levels of noise. Using the measured dependence on the noise, they then extrapolate the result to the expected noiseless value. Instruction-level noise in current digital quantum computers dominantly arises from the two-qubit entangling controlled-NOT (CNOT) gate, and the dominant noise channel is the two-qubit depolarizing channel [31]. In a two-qubit scenario, the depolarizing channel is given by a quantum operation acting on the system’s density matrix $\rho$:
$\mathcal{E}(\rho)=(1-\epsilon)\rho+\frac{\epsilon}{4}I$ (2)
where $\epsilon$ is the error rate parameter and $I$ is the 2 × 2 identity matrix.
We denote the density matrix of a single noisy (depolarizing) CNOT gate on two-qubit $i$ and $j$ by $\rho_{ij}$ and the unitary operator corresponding to the CNOT gate by $U_{c}$. As the action of two CNOT gates is equivalent to the identity operation (see Figure 1 an odd scaling factor $\lambda$ of CNOT gates must be added to the same qubits as follows
$\operatorname{CNOT}_{ij}^{\lambda}[\rho]=(1-\lambda\epsilon)U_{c}\rho U_{c}+\frac{\lambda\epsilon}{4}\rho_{ij}\otimes I_{ij}+O\left(\epsilon^{2}\right)$ (3)
Unitary folding modifies the quantum circuit by incorporating additional operations or ’folds’ that counteract the effects of noise. This strategically introduces corrective operations that enhance the fault-tolerance of the quantum computation, ultimately mitigating the impact of errors.
Parameterized noise scaling counteracts the effects of noise by adjusting the parameters of quantum gates in a controlled manner to account for and counteract the effects of noise. By dynamically scaling the parameters of quantum gates based on the noise characteristics, this technique optimizes the overall performance of the quantum algorithm in noisy environments.
Several methods for scaling quantum circuits have been reported [30, 29, 32, 33]. The two well-known approaches of Giurgica-Tiron et al. [30], i.e., fold from the left and random folding, manipulate noise solely at the gate 0 level to enable ZNE.
  • Noise Scaling from the Left: This approach selects a specific subset of individual gates or layers in a quantum circuit from the left-hand and subject it to noise scaling operations, starting from the leftmost side of the circuit until the end of the quantum circuit. This approach enables targeted noise scaling of specific portions of the circuit, potentially controlling the overall noise characteristics and error mitigation in a controlled manner.
  • Noise Scaling at Random: This approach randomly selects a subset of individual gates or layers for noise scaling operations. Random gate selection enables uniform sampling of the input circuit, smoothing the convergence to a specific scaling factor. By introducing randomness into the gate selection process, this approach can achieve effective noise scaling and error mitigation in QC.
II.3 Noise Extrapolation Models
Extrapolation models for ZNE in quantum computing have been comprehensively discussed elsewhere [22, 30, 10]. The choice of the model depends on the strength of the noise and the number of gates in the circuit. The four different extrapolation models are summarized below.
  • The linear model [22] assumes that the expectation value of a quantum circuit scales linearly with the number of gates. The linear model is simple and can be useful for estimating the zero-noise limit expected value of the circuit when the noise is weak.
  • The exponential model [22] assumes that the noise in a circuit scales exponentially with the number of gates. This model can be used with unitary folding for exponential scaling of the depolarizing parameters of each gate. The obtained parameters are then used for fitting and extrapolating the expectation value of the circuit.
  • The adaptive-exponential model  [30] an extension of the exponential extrapolation model, allows the scaling factor to vary with the number of gates. Estimating the scaling factor using a Bayesian approach, this method improves the accuracy of the extrapolation and reduces the variance of the estimator.
  • The Richardson model [10] maximizes the order of the polynomial extrapolation to the maximum given the number of data points. When the noise is strong and the polynomial order is high, this model improves the accuracy of the extrapolation and reduces the variance of the estimator.
However, unitary folding [30] might not effectively amplify systematic nonuniform noise on the target quantum hardware. Giurgica-Tiron et al. demonstrated that the adaptive-exponential extrapolation method maximizes the accuracy and minimizes the variance, although the exponential, linear, and Richardson extrapolations are useful in different scenarios.
Figure 2: Overall process of using the noise-aware folding. A traditional qubit mapping process is applied to get a mapped circuit. Using the mapped circuit with the target hardware error rate, we can accumulate the noise on the circuit. Finally, we apply the noise-aware folding method to scale the quantum circuit.
III Proposed Noise-aware Folding ZNE
III.1 Overview
Previous methodologies adjusted the gate count of the quantum circuit using the scale factor ($\lambda$), employing techniques such as gate-level unitary folding [30], identity gate insertion [32], and pulse stretching [29] to account for noise scaling. Although these approaches require no specific hardware noise models, they assume uniform noise scaling across the circuit, ignoring the variations in error levels among qubits within the quantum system. In real systems, where the error rates differ at distinct gate levels, these scaling methods cannot obtain uniform scaling factors. Such nonuniform scaling factors can impede the convergence of the extrapolation model toward a zero-noise state effectively.
To tackle these problems, we propose leveraging the noise model of the target hardware to scale the noise on the quantum circuit for ZNE. Our approach regards $\lambda$ not solely as a gate-count scaling factor, but as an error-rate modifier that directly affects the quantum circuit. Quantum computers demand periodic calibration, meaning that calibration data are available for error mitigating by methods such as noise-adaptive compilation [34]. Our approach optimizes the mapping using the noise-adaptive compilation proposed by Murali et al. [34], which favors qubits with high resilience. This method maps logical qubits to physical qubits with low noise levels and minimal distance, thereby reducing the SWAP gate requirements. Employing our noise-aware folding algorithm, we then scale the approximate error rates within the transpiled circuit effectively. The following section explains the specifics of our noise-aware folding method.
III.2 Noise Accumulation
Figure 2 illustrates a quantum circuit and its associated target hardware data, which are input to our proposed method. First, we transpile the quantum circuit using the noise-adaptive compilation mentioned in Section III, efficiently mapping the quantum circuit onto the target hardware by selecting the most robust qubits with low error rates. Although noise is intentionally added to the circuit, noise-adaptive mapping is crucial for ensuring reliable operations, minimizing qubit movement, and reducing the need for SWAP gates, which largely affect the performances of folding methods. To leverage this advantage, a deeper understanding of quantum computer noise—often overlooked in previous ZNE research—is crucial.
The transpiled circuit accumulates noises into an $n$$\times$ $n$ matrix denoted as $qc\_matrix$ in Algorithm 1, where $n$ represents the number of qubits. Along each off-diagonal (where $i$ and $j$ are the rows and column indices, respectively), we aggregate the noise for two-qubit gates between each qubit pair (where $i$ and $j$ denote the indices of the control target qubit, respectively, with $i\neq j$). The noise of the one-qubit gates is accumulated along the diagonal matrix ($i=j$). Because the error rate of a two-qubit gate is independent of the control or target qubit, the noise in the two-qubit gates is accumulated only between qubit pairs located on the upper diagonal of the matrix.
The error rate for each pair of qubits $i$ and $j$, the error rate is aggregated only if these qubits are physically connected and if two-qubit gates exist between them. Once the error rates of the quantum circuit are aggregated into the error rate matrix, the circuit can be scaled at each scale factor.
Input:
$circuit$: input transpiled circuit,
$scale\_factors$: a list of scale factors $\lambda$,
$backend$: target backend quantum computer
Output:
$folded\_circuits$ : a list of scaled circuits
1 Function  fold_noise_aware
2       $cx\_error\_dict$leftarrow GetErrorRateFrom($backend$);
3       $qc\_matrix$leftarrow AccumulateCircuitErrorRate($circuit$);
4       $folded\_circuit$leftarrow empty list;
       // scale the circuit in every $\lambda$ scale factor
5       for $scale$ in $scale\_factors$ do
6             $scaled\_matrix$leftarrow $qc\_matrix$ $scale$;
7             $qc\_rate$leftarrow GetHighestRateIn($qc\_matrix$);
8             $scaled\_rate$leftarrow GetHighestRateIn($scaled\_matrix$) $adjust\_rate$leftarrow ($qc\_rate$ + $scaled\_rate)$ /// $2$;
             // Scale each pair of the qubit on the upper diagonal of the $qc\_matrix$
9             for $i\leftarrow 0$ to $circuit.num\_qubits$ do
10                   for $j\leftarrow$ $i$ to $circuit.num\_qubits$ do
11                         $cur\_rate$leftarrow $qc\_matrix$[i][j];
                         // Perform unitary folding until the adjust rate is reached
12                         while $cur\_rate<adjust\_rate$ do
                               // Get the qubit error rate using control (i) and target (j) qubit as index.
13                               $gate\_rate$leftarrow $cx\_error\_dic[i,j]$;
                               // For every unitary folding insertion, twice the gate $(U)$ error rate is added for $U^{\dagger}U$
14                               $cur\_rate$ += $gate\_rate*2$;
                               // Apply CNOT gate folding to pair i,j
15                               $circuit.cx(i,j)$;
16                               $circuit.cx(i,j)$;
17
18                         end while
19
20                   end for
21
22             end for
23            $folded\_circuit$.insert($circuit$)
24       end for
25      return $folded\_circuit$;
26 end
ALGORITHM 1 Noise-aware Folding
III.3 Noise-aware Folding
As previously discussed, $\lambda$ in our methodology is leveraged to amplify the noise accumulation and is not directly correlated with the circuit’s gate count. This approach involves a nuanced computation (Algorithm 1). By integrating the quantum circuit error rate matrix ($qc\_matrix$ on Line 3) with a scaling constant, we derive the $scaled\_matrix$ on Line 6, effectively delineating the desired error rate augmentation of the scaled $\lambda$ circuit.
To control the pace of noise amplification, our strategy also calculates the average disparity between the maximum error rates within the $scaled\_matrix$ and the ongoing $qc\_matrix$ (Lines 6-9 of Algorithm 1). The computed balanced error rate termed the $adjust\_rate$ on Line 9, is a regulatory mechanism that caps the rate at which errors accumulate.
Next, the error rate is amplified through the unitary folding technique, which substitutes $U$ with $U(U^{\dagger}U)$ to introduce supplementary gates. For every qubit pair within the $qc\_matrix$, this procedure iterates through the gate increments until the existing error rate of the qubit pair ($cur\_rate$) approaches or slightly surpasses the designated $adjust\_rate$ (Lines 10-17 of Algorithm 1).
The method concludes after evaluating all pairs along the upper diagonal matrix. The error rate in resultant circuit is approximately equivalent to the predefined scale $\lambda_{n}=(\lambda_{1}+(\lambda_{1}*\lambda_{n}))/2$.
Importantly, because error rates are available only for gate operations in the noise model, this study exclusively accumulates the noise in two-qubit gates. Nevertheless, our proposed methodology potentially enables seamless extension to all types of gates, provided that adequate error rate calibration data exist within the noise model.
IV Evaluation
This section compares the results of our noise-aware folding method with those of other extrapolation models, including the existing method referenced in [30], using the linear fit extrapolation model.
Figure 3: The circuit configuration designed to benchmark ZNE using various folding methods begins with initializing each qubit to the state $|0\rangle$. Subsequently, a NOT (X) gate is applied to the first qubit, followed by successive CNOT gates between adjacent qubits (Qubit 1 and 2, 2 and 3, …, up to n-1 and n). This sequence results in all qubits being measured in the state $|1\rangle$.
IV.1 Experimental Setup
(a) ibm_mumbai noise model result.
(b) ibm_cairo noise model result.
Figure 4: Comparison of expectation values (vertical axis) obtained through using ZNE linear fit model with different folding methods on a quantum computer simulator, each with its respective noise model. (a) and (b) display results utilizing noise models from ibm_mumbai and ibm_cairo, respectively. The evaluation involves three folding methods: fold from left, fold at random, and noise-aware folding applied to the circuit depicted in Figure 3, spanning from 2 to 27 qubits (horizontal axis). The scaling factor $\lambda$ of $[1,1.5,2,2.5]$ was utilized for each folding method at every qubit count. The ‘Exact Simulation’ represents results without any applied noise model, while the ‘Unmitigated’ results showcase the circuit execution using a noise model without error mitigation. Each result is represented by a distinct symbol and color, as shown in (a).
Our proposed method was validated in various full-noise models from quantum computers available on Qiskit [35]. In these evaluations, we evaluated both the setup circuit in Figure 4 and the Bernstein-Vazirani (BV) circuit as a benchmark [36]. We chose this circuit for benchmarking our proposed method for specific reasons. At a logical level, it presents a relatively straightforward circuit that alters the states of individual qubits. When errors or noise are introduced to this circuit, they are nonnegligibly propagated between qubits and become significant due to their correlation and the necessity to change subsequent qubit states. Any error during execution notably impacts the subsequent qubit within the circuit, depending on the point of its incidence. Moreover, owing to the connectivity constraints, this circuit is not easily executable on quantum computers. Accommodating these limitations in quantum computer connectivity becomes increasingly intricate with the increasing number of qubits.
For each noise model, we ran the quantum circuit with different scaling parameters for comparison purposes. Along with our noise-aware folding method, we employed both from-left and at-random folding methods in [30] implemented on the Mitiq framework [37]. However. owing to the nature of ZNE, we must preserve the folded gates in the folded circuits, which precludes optimization such as gate cancelations before executing the quantum circuit at the compilation level. Given the complexity of our benchmark circuit (Figure 3) and the connectivity limitations of the quantum computer, we must introduce numerous SWAP gates as the number of qubits increases. To avoid this problem and ensure a fair comparison, we preemptively transpile the quantum circuit using Qiskit at the highest $optimization\_level=3$ before initiating the folding process in each method. The resulting transpiled circuit was input to each folding method, each with its own processes for extrapolation.
In addition, all experimental circuits were executed on real quantum computers (see Subsection IV.3). All comparison methods, on both the simulators and real quantum computers, employed the same parameters and compilation process. The subsequent subsection delves into the evaluation results.
IV.2 Result on Simulators
Our experiments were conducted on the currently available 27-qubit system, making use of the full-noise models implemented in Qiskit. Each experiment was conducted five times with different qubit counts and folding methods. Reported are the average outcomes of the five experiments for each folding method. The following graphs present the result of the linear fit extrapolation model, which demonstrated notably reliable results across all qubit counts.
Figure 4 depicts the outcomes of executing the benchmark circuit Figure 3 on three full-noise models, namely, ibm_mumbai and ibm_cairo, using a simulator. As shown in panels (a) and (b) of this figure, the expectation decreased with increasing qubit counts. Applying the unmitigated, fold from left, random folding, and noise-aware folding to the benchmark circuit in the ibm_mumbai noise model, the expectation values respectively decreased from 0.943, 0.952, 0.949, and 0.959 in the 2-qubit circuits to 0.275, 0.347, 0.350, and 0.397 in the 27-qubit circuit (Figure 3(a)).
The expectation values of all execution methods, (including the unmitigated method) were high in the ibm_cairo noise model than in the ibm_mumbai (c.f. Figure 3(a) and  3(b)). This discrepancy stems from differences in the noise model parameters. Specifically, ibm_mumbai yields a significantly higher error rate than the other models.
Figure 5 presents the experimental results of a 14-qubit BV circuit using the full-noise model simulation from ibm_cairo. The expectation value reduced with increasing scaling factor $\lambda$ in our proposed method, but showed no consistent drop in the left and random folding methods. This result demonstrates that in our proposed method, the extrapolation model accurately extrapolates to the zero-noise value.
IV.3 Result on Real Quantum Computers
Figure 6 presents the expectation values on the benchmark circuit in Figure 4, obtained through experiments on a real quantum computer ibm_algiers. The results, which mirror those of Figure 4, were averaged over five executions of each folding method. Notably, the expectation values of the unmitigated, fold from left, fold at random, and noise-aware folding methods declined from 0.979, 0.98, 0.99, and 0.979, respectively, in the 2-qubit circuit to 0.005, 0.021, 0.011, and 0.25, respectively in the 24-qubit circuit. The proposed method consistently outperformed the existing methods from the 2 to 16 qubits and performed comparably to the existing methods beyond the 16-qubit threshold. This discrepancy can be explained by the noticeably faster decline of the expectation value on scaled circuits in our proposed method than in the existing ones. At higher $\lambda$ values, certain folded circuits reach a 0 expectation value, rendering the extrapolated results unreliable.
Comparing executions on real quantum computers to simulations, the overall results deteriorate significantly with increasing qubit counts on real quantum computers. This disparity arises because the noise model does not encompass all noise sources present in real quantum computers during simulations. Consequently, the results are less reliable on real quantum computers.
When the BV circuit was executed on a real quantum computer, the success rate was consistently below 1%. This low success rate translates to an expectation value of less than 0.01 for the unmitigated circuit. Consequently, in folded circuits where the noise is increased, the success rate of the BV circuit approaches 0% and the results become meaningless on this circuit.
Figure 5: Comparison of the expectation value using different folding methods on a Bernstein-Vazirani (BV) [36] circuit with 14-qubit on ibm_cairo noise model simulation. The scale factor shows the drop in expectation value on different folding methods as the circuit gets nosier while $\lambda=0$ is the extrapolated expectation value where the zero-noise is assumed.
Figure 6: Expectation value results on a real quantum computer ibm_algiers with the exact same setup and scaling factor configuration from the result in Figure 4.
V Discussion
We applied other extrapolation methods, namely, exponential fit [22], adaptive-exponential fit from [30], and Richardson’s method [10], to ZNE execution on the above-mentioned circuit folding methods. However, our benchmarking results demonstrated no consistent substantial improvements or more reliable outcomes over those of linear fitting. Indeed, the models were often degraded by extreme overfitting or underfitting, leading to nonconvergence in some instances. Although improvements from those of linear fit extrapolation were sporadically observed, the results of these alternative fittings were too inconsistent for practical execution, particularly in scenarios involving variational-based quantum algorithms.
Notably, the exponential fit and adaptive-exponential fit methods produced identical outputs across all folding methods, including our proposed approach. Therefore, the efficacy and discriminative power of these extrapolations in distinguishing the performances of different extrapolation techniques is questionable. To evaluate our noise-aware folding method against the existing approaches, we conducted comprehensive experiments using linear fit extrapolation. Simulations on various full-noise models demonstrated that our proposed method consistently and efficiently performs across different quantum circuit complexities. Notably, it maintained high extrapolation accuracy as the circuit complexity increased.
Real quantum computer executions also delivered promising results, particularly on smaller circuits. However, at higher qubit counts, the expectation values fell more rapidly with circuit scaling than in the simulation results, affecting the reliability of the extrapolated results. The disparity between the simulation and real quantum computer executions highlights the challenges posed by the noise models, which do not capture all noise sources in actual hardware.
VI Conclusion
Our noise-aware folding method more effectively mitigates quantum errors during computations than the existing methods. By dynamically adjusting the error rates based on hardware noise models, it promises to improve the precision and reliability of quantum computations. The fidelity improvements over the compared methods reached 35% and 26% on quantum computer simulators and real quantum computers, respectively. Although demonstrating robustness in simulations and smaller-scale executions on real quantum computers, the proposed method was challenged by scaling to larger circuits, necessitating further exploration and optimization to broaden its applicability. Owing to its adaptability and versatility, the proposed method can potentially bridge the gap between theoretical expectations and practical outcomes in QC. Further refinements and adaptations are imperative to enhance scalability and reliability, propelling quantum error mitigation methods towards realizing error-free quantum computations.
Acknowledgements.
This research was partly supported by Quantum Computing based on Quantum Advantage challenge research (RS-2023-00257994) through the National Research Foundation of Korea(NRF) funded by the Korean government (MSIT) and Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. 2020-0-00014, A Technology Development of Quantum OS for Fault-tolerant Logical Qubit Computing Environment).

### v2

@@
 Abstract
-The current thousand-qubit processors mark a substantial advance in hardware.
-Yet, hardware limitations prevent quantum error correction (QEC), necessitating reliance on quantum error mitigation (QEM).
-Our paper presents a noise-aware folding method that improves Zero-Noise Extrapolation (ZNE) by estimating noiseless values from noisy results.
-Unlike traditional ZNE methods, which assume a uniform error distribution, our method redistributes the noise using calibration data based on hardware noise models.
-By employing noise-adaptive compilation and optimizing the qubit mappings, our approach enhances the ZNE accuracy of various quantum computing models.
-Recalibrating the noise amplification to address the inherent error variations, promises higher precision and reliability in quantum computations.
-This paper highlights the uniqueness of our method, summarizes noise accumulation, presents the scaling algorithm, and compares the reliability of our method with those of existing models using linear fit extrapolation.
-Relative to the existing folding methods, our method achieved a 35% improvement on quantum computer simulators and a 26% improvement on real quantum computers compared to existing folding methods, demonstrating the effectiveness of our proposed approach.
-preprint: APS/123-QED
-I Introduction
-Recently, quantum computing (QC) technology has entered the era of noisy-intermediate scale quantum computing (NISQ), highlighted by IBM’s recent release of processors housing over a thousand qubits [1, 2]; meanwhile, robust crosstalk chips using tunable-coupler technology, can operate up to 133 qubits [3, 4, 5].
-Despite these advancements, the capacity of current hardware is inadequate for implementing quantum error correction codes but is projected to become adequate by the 2030s [6, 7].
+Recent thousand-qubit processors represent a significant hardware advancement, but current limitations prevent effective quantum error correction (QEC), necessitating reliance on quantum error mitigation (QEM) to enhance result fidelity from quantum computers.
+Our paper introduces a noise-aware folding technique that enhances Zero-Noise Extrapolation (ZNE) by leveraging the noise characteristics of target quantum hardware to fold circuits more efficiently.
+Unlike traditional ZNE approaches assuming uniform error distribution, our method redistributes noise using calibration data based on hardware noise models.
+By employing a noise-adaptive compilation method combined with our proposed folding mechanism, we enhance the ZNE accuracy of quantum gate-based computing using superconducting quantum computers.
+This paper highlights the uniqueness of our method, summarizes noise accumulation, presents the scaling algorithm, and compares the reliability of our method with those of existing models using linear extrapolation model.
+Experimental results show that compared to existing folding methods, our approach achieved a 35% improvement on quantum computer simulators and a 31% improvement on real quantum computers, demonstrating the effectiveness of our proposed approach.
+keywords:
+Quantum computing, quantum error mitigation, zero-noise extrapolation, target-aware optimization
+1 Introduction
+Recently, quantum computing (QC) technology has significant advancements, including IBM’s release of processors housing over a thousand qubits [1, 2], as well as the development of more robust crosstalk chips utilizing tunable-coupler technology with up to 133 qubits [3, 4, 5].
+Despite these advancements, the capacity of current hardware is insufficient for implementing QECs which is projected to become available by the 2030s [6, 7].
 In the meantime, QC systems employ QEM methods, often requiring additional quantum and classical resources as a trade-off to enhance QC output fidelity.
-Numerous QEM techniques for alleviating diverse sources of error in QC have been proposed [8]: probabilistic error cancelation for mitigating decoherence  [9, 10], ZNE for imperfect gates, techniques for mitigating measurement errors  [11, 12, 13], dynamical decoupling [14, 15], quantum optimal control [16, 17], randomized compiling [18, 19], Pauli-frame randomization [20, 21], and other techniques [22, 23, 24, 25, 21, 26].
-The present study focuses on the well-established ZNE technique.
-ZNE was concurrently introduced and extensively demonstrated in various applications involving systems of up to 127 qubits [27, 28].
+Numerous QEM techniques for reducing diverse sources of error in QC have been proposed [8]: probabilistic error cancellation for mitigating decoherence  [9, 10], ZNE for imperfect gates, techniques for mitigating measurement errors  [11, 12, 13], dynamical decoupling [14, 15], quantum optimal control [16, 17], randomized compiling [18, 19], Pauli-frame randomization [20, 21], and other techniques [22, 23, 24, 25, 21, 26].
+Our study focuses on the well-established ZNE technique.
+ZNE was simultaneously introduced and extensively demonstrated in various applications involving systems of up to 127 qubits [27, 28].
 In ZNE, a quantum program is amplified to different noise levels through gate or pulse-level stretching, which intentionally increases the program’s noise.
@@
 When $\lambda=1$, the program operates at its original error rate.
-When $\lambda$ $>0$, an additional error, such as an identity unitary combination of the original gate, is introduced.
-The results obtained at different noise levels of $\lambda$ are collected and extrapolated to $\lambda$ $=0$, effectively eliminating the noise.
-Therefore, $\lambda$ represents the noise level affecting any physical quantity during quantum computation.
+When $\lambda$ $>1$, additional error, such as an identity unitary combination of the original gate, is introduced.
+The results obtained at different noise levels of $\lambda$ are collected and extrapolated to $\lambda$ $=0$, effectively reducing noise to improve the result fidelity.
 Researchers have proposed various methods for amplifying noise in quantum programs.
-For instance, the approach described in [29] stretches gate durations to the desired levels using pulse-level control.
-Under ideal conditions, this stretching does not alter the quantum system’s state, but under noisy conditions, these modifications are classified as errors.
+For instance, the approach described in [29] stretches gate duration to the desired levels using pulse-level control.
+Under ideal conditions, this stretching does not alter the quantum system’s state, but under noisy conditions, these modifications add additional errors.
 Moreover, as this method requires a high degree of abstract control and calibration at the pulse level in quantum computers, it is not easily implementable across most of the existing QC systems.
-In constrast, unitary folding [30] amplifies the noise using gate-level control, which is available in all gate-based computation model QCs.
-This technique builds upon a simple concept: replacing the unitary operation $U\rightarrow U\left(U^{\dagger}U\right)$, where $\left(U^{\dagger}U\right)$ is an identity.
+In contrast, unitary folding [30] amplifies the noise using gate-level control, which is available in all gate-based QCs.
+This technique is based on a simple concept unitary folding: replacing the unitary operation $U\rightarrow U\left(U^{\dagger}U\right)$, where $\left(U^{\dagger}U\right)$ is an identity.
 In an ideal scenario, this replacement remains a $U$ operation, but errors corresponding to U in the quantum system amplify the errors in the $U$ operation within the modified circuit.
 Giurgica-Tiron et al.
-[30] proposed unitary folding with a fold from the left (which folds each unitary gate independently) and at random folding methods (which randomly selects a subset of individual gates randomly as a block of unitary gates and replicates the entire block with $U\left(U^{\dagger}U\right)$.
+[30] proposed unitary folding with a fold from the left (which folds each unitary gate independently) and with random folding methods (which randomly select a subset of individual gates randomly as a block of unitary gates and replicates the entire block with $U\left(U^{\dagger}U\right)$.
 This method requires no knowledge of the QC underlying noise model.
 Although these scaling methods effectively amplify noise in the original input quantum program, they neglect the sources of error imbalance within the QC system, potentially leading to biased results and errors that may compromise the accuracy of the extrapolation results.
-To resolve this problem, we introduce a noise-aware folding method that adjusts the quantum circuit, aiming to balance the error rate across all logical qubits by leveraging the calibration data noise model, which is freely available on select QC systems.
-This approach comprehensively analyzes the calibration data noise model, thereby redistributing the noise amplification process across the quantum circuit.
-Through an algorithmic design, our approach strategically adjusts the scaling factor for each gate operation, balancing the distribution of gate error rates across all logical qubits.
-Addressing the error source imbalances, mitigates bias-induced errors, thereby enhancing the accuracy of ZNE extrapolated results from the quantum program.
-Moreover, the approach is adaptable and therefore versatile across various quantum computing models.
-Seamlessly integrated with gate-based computation models, it comprehensively caters to the inherent complexities of contemporary quantum hardware.
-The integration of a noise-aware folding method enhances the precision and reliability of quantum computations.
-The amplification process can be recalibrated to accommodate the intricate error landscapes within quantum systems, promising more accurate and dependable quantum programming, and bridging the gap between theoretical expectations and practical outcomes.
-The contributions of our study are highlighted below:
-  • Our noise-aware folding method is tailored to redistribute noise across quantum circuits, thereby enhancing the accuracy of ZNE by addressing the inherent error variations within quantum systems.
-  • We seamlessly integrate our approach with gate-based computation models and optimize the qubit mappings using hardware noise models, providing versatility mitigating bias-induced errors, and enhancing the reliability of quantum computations.
-  • Our novel algorithm dynamically adjusts the error rates within the quantum circuit, leveraging calibration data and noise-adaptive compilation methods to ensure a balanced and controlled scaling process.
-  • We conduct rigorous experiments across various full noise models and real quantum computers, showcasing the method’s consistent performance in simulations, highlighting the challenges of scaling to larger circuits, and emphasizing the disparity between simulated and real quantum computer executions.
-Figure 1: Controlled-not (CNOT) operation as identity insertion of unitary U1 and U2 operations using two qubits.
-The remainder of this is organized as follows.
-Section II.1 briefly discusses the background and related QEM methods using ZNE techniques, which motivated our research.
-Section III explains the noise-aware folding method, along with its compilation and execution scheme.
-Section IV analyzes the performance results of our approach and previous methods.
-Our noise-aware folding method is further discussed in Section V and conclusions are presented in Section VI.
-II Background and Related Works
-This section briefly provides the fundamentals of our proposed method. We discuss the quantum error mitigation (QEM), noise scaling methods, and noise extrapolation models.
-II.1 Quantum Error Mitigation
-Quantum computers are susceptible to various sources of noise and imperfections including crosstalk, which can affect the accuracy, reliability, and fidelity of QC results.
-Utilizing a set of techniques and strategies to minimize quantum computational errors, QEM estimates the true expectation value $\langle A\rangle$ from the approximated value $A$ impacted by errors $\psi$.
-$\langle A\rangle=\operatorname{Tr}(A\psi)$ (1)
-QEM strategically minimizes the impact of errors at various stages of quantum processes such as ZNE.
-Ultimately, it aims to optimize the performance of quantum algorithms and facilitate their seamless integration into practical QC environments.
-As one of the best-estimated QC techniques, ZNE is used for understanding the behavior of quantum algorithms in the absence of noise.
-In ZNE, quantum computations are run at different noise levels and the outcomes are extrapolated to a hypothetical noiseless scenario (the zero-noise limit).
-By analyzing the results obtained at various noise levels, we can predict the performance of a quantum algorithm without noise interference.
-This extrapolation provides insights into the intrinsic properties and capabilities of the quantum algorithm, clarifying its idealized behavior.
-Essentially, ZNE can separate the impact of noise from the true potential of a quantum algorithm.
-II.2 Noise Scaling Methods
-ZNE methods measure a given observable at varying levels of noise.
-Using the measured dependence on the noise, they then extrapolate the result to the expected noiseless value.
-Instruction-level noise in current digital quantum computers dominantly arises from the two-qubit entangling controlled-NOT (CNOT) gate, and the dominant noise channel is the two-qubit depolarizing channel [31].
-In a two-qubit scenario, the depolarizing channel is given by a quantum operation acting on the system’s density matrix $\rho$:
-$\mathcal{E}(\rho)=(1-\epsilon)\rho+\frac{\epsilon}{4}I$ (2)
-where $\epsilon$ is the error rate parameter and $I$ is the 2 × 2 identity matrix.
-We denote the density matrix of a single noisy (depolarizing) CNOT gate on two-qubit $i$ and $j$ by $\rho_{ij}$ and the unitary operator corresponding to the CNOT gate by $U_{c}$.
-As the action of two CNOT gates is equivalent to the identity operation (see Figure 1 an odd scaling factor $\lambda$ of CNOT gates must be added to the same qubits as follows
-$\operatorname{CNOT}_{ij}^{\lambda}[\rho]=(1-\lambda\epsilon)U_{c}\rho U_{c}+\frac{\lambda\epsilon}{4}\rho_{ij}\otimes I_{ij}+O\left(\epsilon^{2}\right)$ (3)
-Unitary folding modifies the quantum circuit by incorporating additional operations or ’folds’ that counteract the effects of noise.
-This strategically introduces corrective operations that enhance the fault-tolerance of the quantum computation, ultimately mitigating the impact of errors.
-Parameterized noise scaling counteracts the effects of noise by adjusting the parameters of quantum gates in a controlled manner to account for and counteract the effects of noise.
-By dynamically scaling the parameters of quantum gates based on the noise characteristics, this technique optimizes the overall performance of the quantum algorithm in noisy environments.
-Several methods for scaling quantum circuits have been reported [30, 29, 32, 33].
-The two well-known approaches of Giurgica-Tiron et al.
-[30], i.e., fold from the left and random folding, manipulate noise solely at the gate 0 level to enable ZNE.
-  • Noise Scaling from the Left: This approach selects a specific subset of individual gates or layers in a quantum circuit from the left-hand and subject it to noise scaling operations, starting from the leftmost side of the circuit until the end of the quantum circuit.
-This approach enables targeted noise scaling of specific portions of the circuit, potentially controlling the overall noise characteristics and error mitigation in a controlled manner.
-  • Noise Scaling at Random: This approach randomly selects a subset of individual gates or layers for noise scaling operations.
-Random gate selection enables uniform sampling of the input circuit, smoothing the convergence to a specific scaling factor.
-By introducing randomness into the gate selection process, this approach can achieve effective noise scaling and error mitigation in QC.
-II.3 Noise Extrapolation Models
-Extrapolation models for ZNE in quantum computing have been comprehensively discussed elsewhere [22, 30, 10].
-The choice of the model depends on the strength of the noise and the number of gates in the circuit.
-The four different extrapolation models are summarized below.
-  • The linear model [22] assumes that the expectation value of a quantum circuit scales linearly with the number of gates.
-The linear model is simple and can be useful for estimating the zero-noise limit expected value of the circuit when the noise is weak.
-  • The exponential model [22] assumes that the noise in a circuit scales exponentially with the number of gates.
-This model can be used with unitary folding for exponential scaling of the depolarizing parameters of each gate.
-The obtained parameters are then used for fitting and extrapolating the expectation value of the circuit.
-  • The adaptive-exponential model  [30] an extension of the exponential extrapolation model, allows the scaling factor to vary with the number of gates.
-Estimating the scaling factor using a Bayesian approach, this method improves the accuracy of the extrapolation and reduces the variance of the estimator.
-  • The Richardson model [10] maximizes the order of the polynomial extrapolation to the maximum given the number of data points.
-When the noise is strong and the polynomial order is high, this model improves the accuracy of the extrapolation and reduces the variance of the estimator.
-However, unitary folding [30] might not effectively amplify systematic nonuniform noise on the target quantum hardware.
+To address this issue, we introduce a noise-aware folding method that adjusts the quantum circuit, aiming to balance the error rate across all logical qubits by leveraging the calibration data noise model, which is freely available on most of the QC systems.
+Our approach comprehensively analyzes the calibration data noise model, strategically adjusts the scaling factor for each gate operation, balancing the distribution of gate error rates across all logical qubits.
+Addressing the error source imbalances mitigates bias-induced errors, thereby enhancing the accuracy of ZNE extrapolated results from the quantum program.
+The contributions of our study are highlighted below.
+  • We propose a noise-aware folding method, tailored to redistribute noise across quantum circuits, thereby enhancing the accuracy of ZNE by addressing the inherent error variations within quantum systems.
+  • We seamlessly integrate our approach with gate-based computation models and noise-adaptive compilation using hardware noise models, providing versatility, mitigating bias-induced errors, and enhancing the reliability of quantum computations.
+  • We conduct rigorous experiments across various full-noise models and real quantum computers, showcasing the method’s consistent performance in simulations, highlighting the challenges of scaling to larger circuits, and emphasizing the disparity between simulated and real quantum computer executions.
+The remainder of this paper is organized as follows.
+Section 2 discusses the background and related QEM methods using ZNE techniques, which motivated our research.
+Section 3 explains the noise-aware folding method, along with its compilation and execution scheme.
+Section 4 analyzes the performance results of our approach and previous methods.
+Analytical results are further discussed in Section 5 and conclusions are presented in Section 6.
+2 Background
+This section briefly provides the fundamental concept to understand our proposed method. We present QEM, noise extrapolation model, and noise scaling methods.
+2.1 Zero-noise Extrapolation Method
+Two prominent QEM methods were introduced by Kristan Temme et al.
+[10] in the same letter.
+The first method involves extrapolation to the zero-noise limit (known as ZNE), while the second method is known as Probabilistic Error Cancellation (PEC).
+Both techniques aim to enhance the quality of short-depth quantum circuits and rely on the assumption that a certain noise parameter, combined with system size and circuit depth, can be treated as a small number.
+In terms of noise tolerance, PEC is more robust than ZNE, but it also requires more resources, which increases exponentially with circuit scale.
+As a result, ZNE is more suitable for short-term quantum computers, such as those in the NISQ era.
 Giurgica-Tiron et al.
-demonstrated that the adaptive-exponential extrapolation method maximizes the accuracy and minimizes the variance, although the exponential, linear, and Richardson extrapolations are useful in different scenarios.
-Figure 2: Overall process of using the noise-aware folding.
-A traditional qubit mapping process is applied to get a mapped circuit.
-Using the mapped circuit with the target hardware error rate, we can accumulate the noise on the circuit.
-Finally, we apply the noise-aware folding method to scale the quantum circuit.
-III Proposed Noise-aware Folding ZNE
-III.1 Overview
-Previous methodologies adjusted the gate count of the quantum circuit using the scale factor ($\lambda$), employing techniques such as gate-level unitary folding [30], identity gate insertion [32], and pulse stretching [29] to account for noise scaling.
-Although these approaches require no specific hardware noise models, they assume uniform noise scaling across the circuit, ignoring the variations in error levels among qubits within the quantum system.
-In real systems, where the error rates differ at distinct gate levels, these scaling methods cannot obtain uniform scaling factors.
+[30] outlined the fundamental structure of the ZNE problem in their work.
+The framework assumes that the outcome of a quantum computation is represented by a single expected value denoted as $E(\lambda)$, where $\lambda$ signifies the noise scaling factor.
+This expected value could emerge from either a solitary quantum circuit or a blend of quantum circuits coupled with classical post-processing.
+Notably, $E(\lambda)$ is a real number, ideally estimable only under infinite measurement samples.
+However, in practical scenarios involving a finite number of samples $N$, solely a statistical approximation of the expected value is feasible:
+$\hat{E}(\lambda)=E(\lambda)+\hat{\delta}$ (1)
+Here, $\hat{\delta}$ stands for a stochastic variable with zero mean and variance $\sigma^{2}$.
+While $\hat{E}(\lambda)$ is the estimator for $E(\lambda)$, put it simply $\hat{E}(0)$ for $E(\lambda=0)$ is the estimate of the ZNE with its bias and variance.
+We can draw a genuine prediction $y$ from the probability distribution:
+$P(\hat{E}(\lambda)=y)=N(E(\lambda)-y,\sigma^{2})$ (2)
+In this distribution, $N(\mu,\sigma^{2})$ represents a typical distribution, often Gaussian, characterized by a mean $\mu$ and variance $\sigma^{2}=\sigma_{0}^{2}/N$.
+Given a series of $m$ scaling parameters $\lambda=(\lambda_{1},\lambda_{2},...,\lambda_{m})$, where each $\lambda_{j}$ is greater than or equal to 1, and the corresponding outcomes:
+$y={y_{1},y_{2},...,y_{m}}$ (3)
+Besides the foundational structure ZNE methods, the literature has showcased various models, including linear extrapolation, exponential extrapolation, Richardson extrapolation, among others [30].
+In our paper, we will delve into the specifics of the linear extrapolation model, highlighting its utility and presenting experimental results.
+Linear extrapolation method stands out as one of the simplest techniques, representing a specific instance within the broader scope of polynomial extrapolation.
+This method relies on a polynomial model of degree $d$, expressed as:
+$E_{poly}^{d}=c_{0}+c_{1}\lambda+...c_{d}\lambda^{d}$ (4)
+where $c_{0},c_{1},...,c_{d}$ denote $d+1$ unknown real parameters. From this polynomial model, the linear extrapolation variant emerges:
+$E_{linear}(\lambda)=E_{poly}^{(d=1)}(\lambda)=c_{0}+c_{1}\lambda$ (5)
+Within this context, a straightforward analytical solution exists, manifesting as the ordinary least squared estimator of the intercept parameter:
+$\hat{E}_{linear}(0)=\hat{y}-\frac{S_{\lambda y}}{S_{\lambda\lambda}}\hat{x},$ (6)
+$\hat{\lambda}=\frac{1}{m}\sum_{j}{\lambda_{j}},\hskip 42.67912pt\hat{y}=\frac{1}{m}\sum_{j}{y_{j}},$ (7)
+[…]
+where $\hat{\lambda}$ and $\hat{y}$ represent the averages of the scaling parameters $\lambda_{j}$ and the corresponding outcomes $y_{j}$, respectively.
+The terms $S_{\lambda y}$ and $S_{\lambda\lambda}$ are computed based on the deviations of $\lambda_{j}$ from their mean $\hat{\lambda}$ and the subsequent squares, facilitating the derivation of the intercept parameter.
+This analytical solution underscores the efficiency of the linear extrapolation method, offering a streamlined approach to predict outcomes $y$ with respect to input scaling parameters $\lambda$.
+2.2 Noise Scaling Methods
+Figure 1: Folding methods for ZNE are illustrated with the input circuit depicted in (a), followed by the process of the local unitary folding method to the original circuit in (b), and the global unitary folding method in (c).
+Noticing the difference in the folding process, where $U^{\dagger}U$ was applied: (b) applies each operation to each of the quantum gates individually, while (c) mirrors the entire circuit as a single $U$ operation.
+ZNE method measures a given observable at varying levels of noise.
+Using the measured dependence on the noise, they then extrapolated the result to the expected noiseless value.
+We describe the noise scaling method using two main types of folding namely local unitary folding and global unitary folding:
+  • Unitary folding modifies the quantum circuit by incorporating additional operations or ‘folds’, such as global folding and local folding, which counteract the effects of noise [31].
+This strategically introduces corrective operations that enhance the fault-tolerance quantum computation, ultimately mitigating the impact of errors.
+    $U\rightarrow U\left(U^{\dagger}U\right)^{n},\quad n=0,1,2,\ldots$ (9)
+    In the absence of noise, the replacement rule preserves the operation, as $\left(U^{\dagger}U\right)$ equals the identity.
+However, if noise affects $U$, the unitary folding operation roughly scales the noise by an odd integer factor $\lambda=1+2n$.
+  • Global unitary folding: Assume that the circuit is composed of $d$ unitary layers:
+    $U=L_{d},...,L_{2},L_{1},$ (10)
+    where $d$ represents the depth of the circuit, with each block $L_{j}$ capable of representing a single layer of operations or a solitary gate.
+Within the global folding approach, the substitution rule from Equation 9 extends across the entire circuit, as depicted in Figure 1(c).
+This involves duplicating the entire circuit and adding two additional unitary operations $U$, consolidating all gates within.
+To refine scaling, a final folding can be confined to a subset of the circuit, typically its last $s$ layers.
+Consequently, the general replacement rule for global folding can be summarized as follows.
+    $U\rightarrow U(U^{\dagger}U)^{n}L_{d}^{\dagger}L_{d-1}^{\dagger}...L_{s}^{\dagger}L_{s}...L_{d-1}L_{d}$ (11)
+    The total layer count of the modified circuit is $d(2n+1)+2s$, allowing a stretch of the circuit’s depth to be extended up to a resolution scale of $2/d$.
+This allows for the application of the scaling $d\rightarrow\lambda d$, where
+    $\lambda=1+\frac{2k}{d},k=1,2,3,....$ (12)
+    Conversely, given any real $\lambda$, one can determine the closest integer $k$ to $d(\lambda-1)/2$, perform integer division of $k$ by $d$ to obtain $n$ and $s$, and then apply $n$ integer foldings along with a final partial folding as detailed in Equation 11.
+  • Local unitary folding: Rather than applying global folding to a quantum circuit by appending folds at the end, an alternative approach involves folding a subset of individual gates or layers.
+As depicted in Figure 1(b), each gate undergoes unitary folding individually using Equation 9.
+This results in an evident scaling of the initial number of gates or layers, $d$, by an odd integer, $1+2n$.
+Similar to global folding, a final partial folding operation can be added to achieve a more finely grained scaling factor.
+To implement such ”partial” folding, let’s define an arbitrary subset, $S$, of the full set of indices ${1,2,...,d}$, where the number of elements in $S$ is a specified integer, $s=|S|$.
+Within this framework, the following gate folding rule can be defined:
+    $\forall_{j}\in\{1,2,...,d\},L_{j}\rightarrow\left\{\begin{array}[]{lr}L_{j}(L_{j}^{\dagger}L_{j})^{n},&if~{}j\notin S\\ L_{j}(L_{j}^{\dagger}L_{j})^{n+1},&if~{}j\in S\end{array}\right.$ (13)
+    The selection of elements within the subset $S$ determines the addition of noise channels at various positions throughout the circuit, resulting in various results.
+In the following section, we will explore the expansion of the ”fold from left” rule into a noise-aware folding mechanism.
+This approach involves incorporating error rate data to strategically add layers of gates from the left, deviating from the previously proposed method of folding gates a fixed number of times.
+Figure 2: The entire process of employing noise-aware folding is described in (a), while (b) illustrates the input data, which represents the target quantum hardware.
+Our method begins with the application of a traditional qubit mapping process using the input quantum circuit in (a) to generate a mapped circuit for the target hardware topology in (b).
+Subsequently, the mapped circuit is subjected to the target hardware error rate, which facilitates the accumulation of noise.
+Following this, the noise-aware folding method is implemented.
+Initially, the noise on the circuit is accumulated to form an error rate matrix for $\lambda=1$, representing the original mapped circuit.
+Then, the scaling is applied by computing the scale error rate threshold using the value $\lambda$.
+Finally, identity gates are inserted until the error rate of the qubit pair approaches the threshold.
+3 Proposed Noise-aware Folding ZNE
+3.1 Overview
+As previously outlined, we adjusted the gate count of the quantum circuit using the scale factor ($\lambda$) through gate-level unitary folding.
+However, additional techniques such as identity gate insertion [32] and pulse stretching [29] are also utilized to accommodate noise scaling.
+Although some of these approaches require no specific hardware noise models, they assume uniform noise scaling across the circuit, ignoring the variations in error levels among qubits within the quantum system.
+In real systems, where error rates differ at different logical gates for each physical qubit, these scaling methods cannot obtain uniform scaling factors.
 Such nonuniform scaling factors can impede the convergence of the extrapolation model toward a zero-noise state effectively.
-To tackle these problems, we propose leveraging the noise model of the target hardware to scale the noise on the quantum circuit for ZNE.
-Our approach regards $\lambda$ not solely as a gate-count scaling factor, but as an error-rate modifier that directly affects the quantum circuit.
-Quantum computers demand periodic calibration, meaning that calibration data are available for error mitigating by methods such as noise-adaptive compilation [34].
+To address these problems, we propose to leverage the noise model of the target hardware to scale the noise using a noise-aware folding method in ZNE for the quantum circuit.
+Our approach regards $\lambda$ not solely as a gate-count scaling factor which is the main differ from fold from left method, but as an error-rate modifier threshold to scale the quantum circuit.
+Quantum computers require periodic calibration, which means that calibration data are readily available to the public.
+Although the calibration data is limited to a certain logical gates such as (one- and two-qubit gate) especially the native gate of the quantum computer, our proposed method can be used with the available data as little as a calibration error rate of a single logical two-qubit gate.
 Our approach optimizes the mapping using the noise-adaptive compilation proposed by Murali et al.
 [33], which favors qubits with high resilience.
-This method maps logical qubits to physical qubits with low noise levels and minimal distance, thereby reducing the SWAP gate requirements.
-Employing our noise-aware folding algorithm, we then scale the approximate error rates within the transpiled circuit effectively.
+This method maps logical qubits to physical qubits with low noise levels and minimal distance, thus reducing the SWAP gate requirements.
+Employing our noise-aware folding algorithm, we then effectively scale the approximate error rates within the mapped circuit.
 The following section explains the specifics of our noise-aware folding method.
-III.2 Noise Accumulation
-Figure 2 illustrates a quantum circuit and its associated target hardware data, which are input to our proposed method.
-First, we transpile the quantum circuit using the noise-adaptive compilation mentioned in Section III, efficiently mapping the quantum circuit onto the target hardware by selecting the most robust qubits with low error rates.
-Although noise is intentionally added to the circuit, noise-adaptive mapping is crucial for ensuring reliable operations, minimizing qubit movement, and reducing the need for SWAP gates, which largely affect the performances of folding methods.
-To leverage this advantage, a deeper understanding of quantum computer noise—often overlooked in previous ZNE research—is crucial.
-The transpiled circuit accumulates noises into an $n$$\times$ $n$ matrix denoted as $qc\_matrix$ in Algorithm 1, where $n$ represents the number of qubits.
-Along each off-diagonal (where $i$ and $j$ are the rows and column indices, respectively), we aggregate the noise for two-qubit gates between each qubit pair (where $i$ and $j$ denote the indices of the control target qubit, respectively, with $i\neq j$).
-The noise of the one-qubit gates is accumulated along the diagonal matrix ($i=j$).
-Because the error rate of a two-qubit gate is independent of the control or target qubit, the noise in the two-qubit gates is accumulated only between qubit pairs located on the upper diagonal of the matrix.
+3.2 Noise Accumulation
+Figure 2 depicts an input quantum circuit alongside the target quantum hardware properties, encompassing the qubit connectivity topology and the calibration data for 2-qubit error rates.
+In this illustration, we focus on the two-qubit gate due to its higher error rate compared to a one-qubit gate.
+Nevertheless, our method is readily adaptable to support both one- and two-qubit gates.
+This visualization helps to understand how the circuit will be mapped onto the hardware and how errors may affect its execution.
+The error rate matrix contains a 2-qubit error rate for all of the qubits that have a connection to another qubit.
+First, the input quantum circuit undergoes noise-adaptive compilation, as described in Section 3, efficiently mapping the quantum circuit onto the target hardware by selecting the most robust qubits with low error rates, shown in Figure 2.
+In this process, logical qubits $Q_{1}$, $Q_{2}$, and $Q_{3}$ are mapped to physical qubits $1$, 00, and $2$, respectively.
+These specific physical qubits were chosen because their error rates for 2-qubit gates are lower compared to the rest of the qubits in the system.
+Although noise is intentionally added to the circuit in the next process of all of the folding methods, noise-adaptive mapping is crucial for ensuring reliable operations, minimizing qubit movement, and reducing the need for SWAP gates, which largely affect the performances of folding methods.
+To leverage this advantage, a deeper understanding of quantum computer noise—often overlooked in previous ZNE researches is crucial.
+Let’s consider a mapped circuit comprised of $d$ unitary layers.
+However, unlike Equation 10 where each layer $L$ represents a unitary gate, here $L$ can be either a one- or two-qubit gate.
+Thus, the circuit can be represented as
+$U=L_{d}...L_{2}L_{1},$ (14)
+where for each $j\in{1,2,...,d}$, $L_{j}$ is defined as:
+$L_{j}\rightarrow\left\{\begin{array}[]{lr}L_{j}\rightarrow g_{i,k},&if~{}L_{j}\textit{is a two-qubit gate}\\ L_{j}\rightarrow g_{i},&if~{}L_{j}\textit{is a one-qubit gate}\end{array}\right.$ (15)
+Here, $i$ and $k$ represent logical qubits, where $g_{i,k}$ denotes a two-qubit gate acting on the control qubit $i$ and the target qubit $k$, and $g_{i}$ represents a one-qubit gate acting solely on the target qubit $i$.
+The mapped circuit accumulates noises into an $n$$\times$ $n$ matrix denoted as $qc\_matrix$ in Algorithm 1, where $n$ represents the number of qubits.
+In the matrix, the number of rows and columns represents the physical qubit.
+Along each off-diagonal (where $i$ and $k$ are the rows and column indices, respectively), we aggregate the noise for two-qubit gates $g_{i,k}$ between each qubit pair (where $i$ and $k$ denote the indices of the control target qubit, respectively, with $i\neq k$) by simply summing the total error rate value using the input calibration error rate to the number of two-qubit gate $g_{i,k}$.
+The noise accumulation in this phase represents $\lambda=1$, indicating that the folding methods do not introduce any additional gates or errors to the circuit.
 The error rate for each pair of qubits $i$ and $j$, the error rate is aggregated only if these qubits are physically connected and if two-qubit gates exist between them.
+While along each diagonal in the $qc\_matrix$, one-qubit gate $g_{i}$ are accumulated.
+However in our example, in Figure 2 there are no one qubit-gate, thus making the value along the diagonal all 00.
 Once the error rates of the quantum circuit are aggregated into the error rate matrix, the circuit can be scaled at each scale factor.
+Subsequent noise-aware folding processes for $\lambda$ are discussed in the next section.
 Input:
@@
 $folded\_circuits$ : a list of scaled circuits
-1 Function  fold_noise_aware
+1
+2Function  fold_noise_aware
+3
 4      $cx\_error\_dict$leftarrow GetErrorRateFrom($backend$);
@@
 6      $folded\_circuit$leftarrow empty list;
-       // scale the circuit in every $\lambda$ scale factor
+      /* scale the circuit in every $\lambda$ scale factor */
 7       for $scale$ in $scale\_factors$ do
+8
 9            $scaled\_matrix$leftarrow $qc\_matrix$ $scale$;
 10            $qc\_rate$leftarrow GetHighestRateIn($qc\_matrix$);
-8             $scaled\_rate$leftarrow GetHighestRateIn($scaled\_matrix$) $adjust\_rate$leftarrow ($qc\_rate$ + $scaled\_rate)$ /// $2$;
-             // Scale each pair of the qubit on the upper diagonal of the $qc\_matrix$
+11            $scaled\_rate$leftarrow GetHighestRateIn($scaled\_matrix$)
+12            $adjust\_rate$leftarrow ($qc\_rate$ + $scaled\_rate)$ /// $2$;
+            /* Scale each pair of the qubit on the upper diagonal of the $qc\_matrix$/
 13             for $i\leftarrow 0$ to $circuit.num\_qubits$ do
+14
 15                  for $j\leftarrow$ $i$ to $circuit.num\_qubits$ do
+16
 17                        $cur\_rate$leftarrow $qc\_matrix$[i][j];
-                         // Perform unitary folding until the adjust rate is reached
+                        /* Perform unitary folding until the adjust rate is reached */
 18                         while $cur\_rate<adjust\_rate$ do
-                               // Get the qubit error rate using control (i) and target (j) qubit as index.
+19
+                              /* Get the qubit error rate using control (i) and target (j) qubit as index. */
 20                               $gate\_rate$leftarrow $cx\_error\_dic[i,j]$;
-                               // For every unitary folding insertion, twice the gate $(U)$ error rate is added for $U^{\dagger}U$
+                              /* For every unitary folding insertion, twice the gate $(U)$ error rate is added for $U^{\dagger}U$/
 21                               $cur\_rate$ += $gate\_rate*2$;
-                               // Apply CNOT gate folding to pair i,j
+                              /* Apply CNOT gate folding to pair i,j */
 22                               $circuit.cx(i,j)$;
 23                              $circuit.cx(i,j)$;
-17
 24                         end while
-19
+25
 26                   end for
-21
+27
 28             end for
@@
 30       end for
+31
 32      return $folded\_circuit$;
@@
 ALGORITHM 1 Noise-aware Folding
-III.3 Noise-aware Folding
-As previously discussed, $\lambda$ in our methodology is leveraged to amplify the noise accumulation and is not directly correlated with the circuit’s gate count.
-This approach involves a nuanced computation (Algorithm 1).
-By integrating the quantum circuit error rate matrix ($qc\_matrix$ on Line 3) with a scaling constant, we derive the $scaled\_matrix$ on Line 6, effectively delineating the desired error rate augmentation of the scaled $\lambda$ circuit.
-To control the pace of noise amplification, our strategy also calculates the average disparity between the maximum error rates within the $scaled\_matrix$ and the ongoing $qc\_matrix$ (Lines 6-9 of Algorithm 1).
-The computed balanced error rate termed the $adjust\_rate$ on Line 9, is a regulatory mechanism that caps the rate at which errors accumulate.
-Next, the error rate is amplified through the unitary folding technique, which substitutes $U$ with $U(U^{\dagger}U)$ to introduce supplementary gates.
-For every qubit pair within the $qc\_matrix$, this procedure iterates through the gate increments until the existing error rate of the qubit pair ($cur\_rate$) approaches or slightly surpasses the designated $adjust\_rate$ (Lines 10-17 of Algorithm 1).
-The method concludes after evaluating all pairs along the upper diagonal matrix.
-The error rate in resultant circuit is approximately equivalent to the predefined scale $\lambda_{n}=(\lambda_{1}+(\lambda_{1}*\lambda_{n}))/2$.
-Importantly, because error rates are available only for gate operations in the noise model, this study exclusively accumulates the noise in two-qubit gates.
-Nevertheless, our proposed methodology potentially enables seamless extension to all types of gates, provided that adequate error rate calibration data exist within the noise model.
-IV Evaluation
+3.3 Noise-aware Folding
+As previously discussed, $\lambda$ in our methodology is leveraged to amplify the noise accumulation and is not directly correlated with the circuit’s gate count or depth as previously discussed on global or local unitary folding.
+To achieve this, we propose a noise scaling based on the error rate with the adjustment using $\lambda$ as follow:
+$\epsilon_{max}=(\epsilon_{circuit}+\epsilon_{\lambda})/\gamma,$ (16)
+where $\epsilon_{circuit}$ is a real number representing the highest error rate in the matrix accumulated at $\lambda=1$, as shown in Figure 2.
+Here, $\epsilon_{\lambda}$ is the error rate adjusted by $\lambda$, calculated as $\epsilon_{\lambda}=\epsilon_{circuit}*\lambda$.
+Finally, $\gamma$ is an adjustable coefficient parameter that allows control the rate of the scaling.
+In Figure 2 and algorithm 1, $\gamma$ is set to $2$.
+This computation is detailed in the algorithm 1 line 6-9.
+Using $\epsilon_{max}$ at any $\lambda$, our noise-aware folding mechanism applies the unitary folding using Equation 9 until the folded gate error accumulation reaches close or equal to $\epsilon_{max}$, as detailed in algorithm 1 line 13-17.
+This approach is noise-aware and significantly differs from previous methods because the rate at which unitary folding is applied to the circuit depends entirely on the error rate of the physical qubit.
+As illustrated in Figure 2, at $\lambda=2$, unitary folding was applied equally by adding 2 CNOT gates to pair of logical qubit $Q_{1,2}$ and $Q_{2,3}$ on the quantum circuit.
+However notice at $\lambda=4$, 4 CNOT gates were added to the pair of logical qubit $Q_{1,2}$ while only 2 CNOT gates were added to the pair of logical qubit $Q_{2,3}$ in the quantum circuit despite pair of $Q_{1,2}$ has more CNOT gates and a SWAP gate as well.
+This is because the accumulated error rate on the pair of qubit $Q_{1,2}$ and $Q_{2,3}$ are 0.063 and 0.057, respectively, with the $\epsilon_{max}=0.063$.
+$Q_{1,2}$ reaches the maximum threshold, while $Q_{2,3}$ reaches close to maximum and cannot add additional $U^{\dagger}U$ anymore due to it would be exceed $\epsilon_{max}$ and such condition is not allowed.
+Note that this study focuses solely on applying the folding method for two-qubit gates due to their notable error rates and the availability of calibration data.
+However, our proposed noise-aware folding method can be expanded to accommodate all gate types, assuming sufficient error rate calibration data are available within the noise model.
+4 Evaluation
 This section compares the results of our noise-aware folding method with those of other extrapolation models, including the existing method referenced in [30], using the linear fit extrapolation model.
@@
 This sequence results in all qubits being measured in the state $|1\rangle$.
-IV.1 Experimental Setup
+4.1 Experimental Setup
 (a) ibm_mumbai noise model result.
@@
 Figure 4: Comparison of expectation values (vertical axis) obtained through using ZNE linear fit model with different folding methods on a quantum computer simulator, each with its respective noise model.
-(a) and (b) display results utilizing noise models from ibm_mumbai and ibm_cairo, respectively.
+(a) and (b) show results using noise models from ibm_mumbai and ibm_cairo, respectively.
 The evaluation involves three folding methods: fold from left, fold at random, and noise-aware folding applied to the circuit depicted in Figure 3, spanning from 2 to 27 qubits (horizontal axis).
@@
 In these evaluations, we evaluated both the setup circuit in Figure 3 and the Bernstein-Vazirani (BV) circuit as a benchmark [35].
-We chose this circuit for benchmarking our proposed method for specific reasons.
+We selected this circuit (see Figure 3) to assess our proposed method due to its similarity to the fully-entangled layer utilized in various NISQ applications.
 At a logical level, it presents a relatively straightforward circuit that alters the states of individual qubits.
-When errors or noise are introduced to this circuit, they are nonnegligibly propagated between qubits and become significant due to their correlation and the necessity to change subsequent qubit states.
+When errors or noise are introduced to this circuit, they are non-negligibly propagated between qubits and become significant due to their correlation and the necessity to change subsequent qubit states.
 Any error during execution notably impacts the subsequent qubit within the circuit, depending on the point of its incidence.
 Moreover, owing to the connectivity constraints, this circuit is not easily executable on quantum computers.
-Accommodating these limitations in quantum computer connectivity becomes increasingly intricate with the increasing number of qubits.
+Accommodating these limitations in quantum computer connectivity becomes increasingly complex with the increasing number of qubits.
 For each noise model, we ran the quantum circuit with different scaling parameters for comparison purposes.
 Along with our noise-aware folding method, we employed both from-left and at-random folding methods in [30] implemented on the Mitiq framework [36].
-However.
-owing to the nature of ZNE, we must preserve the folded gates in the folded circuits, which precludes optimization such as gate cancelations before executing the quantum circuit at the compilation level.
+However, owing to the nature of ZNE, we must preserve the folded gates in the folded circuits, which precludes optimization such as gate cancelations before executing the quantum circuit at the compilation level.
 Given the complexity of our benchmark circuit (Figure 3) and the connectivity limitations of the quantum computer, we must introduce numerous SWAP gates as the number of qubits increases.
@@
 The resulting transpiled circuit was input to each folding method, each with its own processes for extrapolation.
-In addition, all experimental circuits were executed on real quantum computers (see Subsection IV.3).
-All comparison methods, on both the simulators and real quantum computers, employed the same parameters and compilation process.
+In addition, all experimental circuits were executed on real quantum computers (see Subsection 4.3).
+All comparison methods, both on simulators and on real quantum computers, employed the same parameters and compilation process.
 The subsequent subsection delves into the evaluation results.
-IV.2 Result on Simulators
-Our experiments were conducted on the currently available 27-qubit system, making use of the full-noise models implemented in Qiskit.
+Figure 5: Comparison of the expectation value using different folding methods on a Bernstein-Vazirani (BV) [35] circuit with 14-qubit on ibm_cairo noise model simulation.
+The scale factor shows the drop in expectation value in different folding methods as the circuit gets nosier, while $\lambda=0$ is the extrapolated expectation value where zero noise is assumed.
+4.2 Result on Simulators
+Our experiments were conducted on the 27-qubit system, making use of the full-noise models implemented in Qiskit.
 Each experiment was conducted five times with different qubit counts and folding methods.
-Reported are the average outcomes of the five experiments for each folding method.
+The results are the average of the five experiments for each folding method.
 The following graphs present the result of the linear fit extrapolation model, which demonstrated notably reliable results across all qubit counts.
-Figure 4 depicts the outcomes of executing the benchmark circuit Figure 3 on three full-noise models, namely, ibm_mumbai and ibm_cairo, using a simulator.
+Figure 4 depicts the results of executing the benchmark circuit shown in Figure 3 on two full-noise models, namely, ibm_mumbai and ibm_cairo, using a simulator.
 As shown in panels (a) and (b) of this figure, the expectation decreased with increasing qubit counts.
-Applying the unmitigated, fold from left, random folding, and noise-aware folding to the benchmark circuit in the ibm_mumbai noise model, the expectation values respectively decreased from 0.943, 0.952, 0.949, and 0.959 in the 2-qubit circuits to 0.275, 0.347, 0.350, and 0.397 in the 27-qubit circuit (Figure 3(a)).
+Applying the unmitigated, fold from left, random folding, and noise-aware folding to the benchmark circuit in the ibm_mumbai noise model, the expectation values respectively decreased from 0.943, 0.952, 0.949, and 0.959 in the 2-qubit circuits to 0.237, 0.292, 0.279, and 0.318 in the 27-qubit circuit (Figure 4(a)).
 The expectation values of all execution methods, (including the unmitigated method) were high in the ibm_cairo noise model than in the ibm_mumbai (c.f.
-Figure 3(a) and  3(b)).
-This discrepancy stems from differences in the noise model parameters.
+Figure 4(a) and  4(b)).
+This discrepancy arises from differences in the noise model parameters.
 Specifically, ibm_mumbai yields a significantly higher error rate than the other models.
 Figure 5 presents the experimental results of a 14-qubit BV circuit using the full-noise model simulation from ibm_cairo.
-The expectation value reduced with increasing scaling factor $\lambda$ in our proposed method, but showed no consistent drop in the left and random folding methods.
+The expectation value reduced with increasing scaling factor $\lambda$ in our proposed method, but did not show a consistent drop in the left and random folding methods.
 This result demonstrates that in our proposed method, the extrapolation model accurately extrapolates to the zero-noise value.
-IV.3 Result on Real Quantum Computers
-Figure 6 presents the expectation values on the benchmark circuit in Figure 4, obtained through experiments on a real quantum computer ibm_algiers.
+Figure 6: Expectation value results on a real quantum computer ibmq_mumbai with the exact same setup and scaling factor configuration from the result in Figure 4.
+4.3 Result on Real Quantum Computer
+Figure 6 presents the expectation values on the benchmark circuit shown in Figure 3, obtained through experiments on a real quantum computer ibmq_mumbai.
+As of preparing this manuscript, IBM has retired ibmq_mumbai; however, we captured its noise model at exact same time that we performed the experiment.
+You can find details regarding its properties and noise model in the appendix.
 The results, which mirror those of Figure 4, were averaged over five executions of each folding method.
-Notably, the expectation values of the unmitigated, fold from left, fold at random, and noise-aware folding methods declined from 0.979, 0.98, 0.99, and 0.979, respectively, in the 2-qubit circuit to 0.005, 0.021, 0.011, and 0.25, respectively in the 24-qubit circuit.
-The proposed method consistently outperformed the existing methods from the 2 to 16 qubits and performed comparably to the existing methods beyond the 16-qubit threshold.
+Notably, the expectation values of the unmitigated, fold from left, fold at random, and noise-aware folding methods declined from 0.949, 0.956, 0.953, and 0.981, respectively, in the 2-qubit circuit to 0.007, 0.019, 0.015, and 0.12, respectively in the 24-qubit circuit.
+The proposed method consistently outperformed the existing methods from the 2 to 21 qubits.
 This discrepancy can be explained by the noticeably faster decline of the expectation value on scaled circuits in our proposed method than in the existing ones.
-At higher $\lambda$ values, certain folded circuits reach a 0 expectation value, rendering the extrapolated results unreliable.
-Comparing executions on real quantum computers to simulations, the overall results deteriorate significantly with increasing qubit counts on real quantum computers.
-This disparity arises because the noise model does not encompass all noise sources present in real quantum computers during simulations.
-Consequently, the results are less reliable on real quantum computers.
+At higher $\lambda$ values with those of higher than 22-qubit, certain folded circuits reach a 0 expectation value, rendering the extrapolated results unreliable.
+Comparing executions on real quantum computers to simulations, the overall results drop significantly with increasing qubit counts on real quantum computers.
+This disparity arises because the noise model cannot fully represent all noise sources in real quantum computers during simulations making the results on real quantum computers less reliable.
 When the BV circuit was executed on a real quantum computer, the success rate was consistently below 1%.
@@
 Consequently, in folded circuits where the noise is increased, the success rate of the BV circuit approaches 0% and the results become meaningless on this circuit.
-Figure 5: Comparison of the expectation value using different folding methods on a Bernstein-Vazirani (BV) [36] circuit with 14-qubit on ibm_cairo noise model simulation.
-The scale factor shows the drop in expectation value on different folding methods as the circuit gets nosier while $\lambda=0$ is the extrapolated expectation value where the zero-noise is assumed.
-Figure 6: Expectation value results on a real quantum computer ibm_algiers with the exact same setup and scaling factor configuration from the result in Figure 4.
-V Discussion
+5 Discussion
 We applied other extrapolation methods, namely, exponential fit [22], adaptive-exponential fit from [30], and Richardson’s method [10], to ZNE execution on the above-mentioned circuit folding methods.
-However, our benchmarking results demonstrated no consistent substantial improvements or more reliable outcomes over those of linear fitting.
-Indeed, the models were often degraded by extreme overfitting or underfitting, leading to nonconvergence in some instances.
-Although improvements from those of linear fit extrapolation were sporadically observed, the results of these alternative fittings were too inconsistent for practical execution, particularly in scenarios involving variational-based quantum algorithms.
-Notably, the exponential fit and adaptive-exponential fit methods produced identical outputs across all folding methods, including our proposed approach.
-Therefore, the efficacy and discriminative power of these extrapolations in distinguishing the performances of different extrapolation techniques is questionable.
-To evaluate our noise-aware folding method against the existing approaches, we conducted comprehensive experiments using linear fit extrapolation.
-Simulations on various full-noise models demonstrated that our proposed method consistently and efficiently performs across different quantum circuit complexities.
-Notably, it maintained high extrapolation accuracy as the circuit complexity increased.
-Real quantum computer executions also delivered promising results, particularly on smaller circuits.
-However, at higher qubit counts, the expectation values fell more rapidly with circuit scaling than in the simulation results, affecting the reliability of the extrapolated results.
-The disparity between the simulation and real quantum computer executions highlights the challenges posed by the noise models, which do not capture all noise sources in actual hardware.
-VI Conclusion
-Our noise-aware folding method more effectively mitigates quantum errors during computations than the existing methods.
-By dynamically adjusting the error rates based on hardware noise models, it promises to improve the precision and reliability of quantum computations.
-The fidelity improvements over the compared methods reached 35% and 26% on quantum computer simulators and real quantum computers, respectively.
+However, the results demonstrated no consistent substantial improvements or more reliable outcomes over those of linear fitting model.
+In fact, the models were often degraded by extreme over-fitting or under-fitting, leading to non-convergence in some instances.
+Although improvements from those of linear fit extrapolation were sometimes observed, the results of these fittings were too inconsistent for practical execution, particularly in scenarios involving variational-based quantum algorithms where a consistent result is required for many iterations.
+Notably, the exponential fit and adaptive-exponential fit methods produced identical outputs across all folding methods, including our proposed noise-aware folding approach.
+Hence, the effectiveness of these extrapolations in discerning the performances of various extrapolation techniques are uncertain.
+Although linear fit extrapolation, like any other models, is unable to extrapolate to zero noise as theoretically proposed due to strong noise in the quantum computer, the consistent improvement in results we observe using linear fit sets it apart from other extrapolation models.
+However, utilizing zero-noise extrapolation comes with a trade-off of executing more quantum circuits, typically ranging from $2x$ to $5x$ the original amount, depending on the number of $\lambda$ values desired by the user.
+Moreover, while QEM techniques like ZNE are primarily employed to enhance the results of NISQ devices, they may also prove practical for use in quantum error correction codes to protect qubits from errors.
+By employing both error correction codes and mitigation strategies, a more resource-efficient method can be achieved compared to traditional error correction codes alone.
+In this hybrid approach, QEM techniques can be used to pre-process the noisy qubits before encoding them into a QEC code.
+This can reduce the error rates of the qubits, making the subsequent QEC code more effective and efficient.
+For instance, ZNE can be applied to suppress errors in the qubits before encoding them into a surface code or a concatenated code.
+Alternatively, QEM techniques can be integrated directly into the QEC code, allowing for real-time error mitigation during the execution of the quantum algorithm.
+This can be particularly useful for codes that are prone to error accumulation, such as those used in quantum simulations or machine learning applications.
+As advancements in quantum computer technology reduce noise levels, making error mitigation strategies more effective, this potential becomes even more viable.
+With lower noise levels, QEM techniques can provide a greater reduction in error rates, further enhancing the performance of the hybrid approach.
+6 Conclusion
+We proposed a noise-aware folding method to improve the zero-noise extrapolation method by employing the noise-adaptive compilation and a folding mechanism that distributes noise uniformly using calibration data based on the quantum hardware noise model.
+By dynamically adjusting the error rates based on hardware noise models, it improves the precision and reliability of the linear fit extrapolation method results in improving the quantum computer result fidelity The fidelity improvements over the compared methods reached 35% and 31% on quantum computer simulators and real quantum computers, respectively.
 Although demonstrating robustness in simulations and smaller-scale executions on real quantum computers, the proposed method was challenged by scaling to larger circuits, necessitating further exploration and optimization to broaden its applicability.
-Owing to its adaptability and versatility, the proposed method can potentially bridge the gap between theoretical expectations and practical outcomes in QC.
-Further refinements and adaptations are imperative to enhance scalability and reliability, propelling quantum error mitigation methods towards realizing error-free quantum computations.
-Acknowledgements.
-This research was partly supported by Quantum Computing based on Quantum Advantage challenge research (RS-2023-00257994) through the National Research Foundation of Korea(NRF) funded by the Korean government (MSIT) and Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.
+Further refinements and adaptations are essential to enhance scalability and reliability, driving quantum error mitigation methods towards realizing error-free quantum computations.
+[…]
+Acknowledgments This research was partly supported by Quantum Computing based on Quantum Advantage challenge research (RS-2023-00257994) through the National Research Foundation of Korea(NRF) funded by the Korean government (MSIT) and Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.
 2020-0-00014, A Technology Development of Quantum OS for Fault-tolerant Logical Qubit Computing Environment).
+Appendix A ibmq_mumbai’s Properties
+This supplementary document provides the experimental setup detail of the quantum computer that we use in our demonstration.
+Figure 7: ibmq_mumbai’s qubit connectivity topology.
+Qubit T1 (us) T2 (us) Frequency (GHz) Anharmonicity (GHz)
+0 111.6961464 179.1588076 5.070549646 -0.328451763
+1 122.2484892 197.5837264 4.929863389 -0.331291977
+2 81.90697735 155.8248129 4.670281392 -0.336876757
+3 167.7648109 122.141877 4.889402396 -0.331179264
+4 42.05229013 29.68056521 5.020938057 -0.330174519
+5 101.1535787 122.3569753 4.969532374 -0.329757062
+6 103.291911 45.33644272 4.965898301 -0.329305699
+7 67.91110372 48.58658623 4.894369882 -0.331001624
+8 87.45290954 142.5387495 4.79164979 -0.332559138
+9 96.66829381 73.24259076 4.954976802 -0.330638393
+10 110.9073585 132.0053088 4.958620128 -0.330907906
+11 88.46621356 202.3471754 4.666070269 -0.332629168
+12 141.6272911 204.5251318 4.74291957 -0.332978886
+13 117.1107884 151.0626282 4.88858331 -0.328060893
+14 84.32426156 153.5437913 4.780232309 -0.332554813
+15 97.45073523 90.18823679 4.8581701 -0.333242963
+16 104.390153 262.8877462 4.98035969 -0.329920963
+17 104.1706916 85.0072165 5.003339003 -0.329925892
+18 122.1904549 212.2866827 4.781291798 -0.333097806
+19 117.3925825 251.3554841 4.809727654 -0.33209643
+20 130.2027001 214.4300921 5.047734483 -0.328048366
+21 109.9161771 200.7424141 4.942677405 -0.331299817
+22 130.3648863 211.195153 4.911344811 -0.331838149
+23 90.39455816 163.0951773 4.892629733 -0.331514799
+24 141.7988532 29.62637317 4.671041833 -0.335931163
+25 148.610822 128.6915538 4.758538084 -0.333550575
+26 72.49759913 135.6307972 4.954268232 -0.329535731
+Table 1: ibmq_mumber’s noise model captured on 26-March-2024 from IBM.
+Qubit Readout assignment error Prob meas0 prep1 Prob meas1 prep0 Readout length (ns) ID error Sx error Pauli-X error CNOT error
+0 2.26E-02 3.54E-02 9.80E-03 3512.888889 2.74E-04 2.74E-04 2.74E-04 0_1:5.62E-03
+1 1.58E-02 2.62E-02 5.40E-03 3512.888889 1.93E-04 1.93E-04 1.93E-04 1_2:8.25E-03;1_4:5.29E-03;1_0:5.62E-03
+2 1.15E-02 1.60E-02 7.00E-03 3512.888889 1.16E-04 1.16E-04 1.16E-04 2_3:5.11E-03;2_1:8.25E-03
+3 1.15E-02 1.74E-02 5.60E-03 3512.888889 2.33E-04 2.33E-04 2.33E-04 3_2:5.11E-03;3_5:7.57E-03
+4 2.73E-02 3.72E-02 1.74E-02 3512.888889 2.77E-04 2.77E-04 2.77E-04 4_7:8.49E-03;4_1:5.29E-03
+5 1.74E-02 2.32E-02 1.16E-02 3512.888889 2.33E-04 2.33E-04 2.33E-04 5_3:7.57E-03;5_8:2.25E-02
+6 1.72E-02 2.58E-02 8.60E-03 3512.888889 2.57E-04 2.57E-04 2.57E-04 6_7:1.22E-02
+7 1.88E-02 3.12E-02 6.40E-03 3512.888889 4.94E-04 4.94E-04 4.94E-04 7_10:8.76E-03;7_6:1.22E-02;7_4:8.49E-03
+8 1.97E-02 2.96E-02 9.80E-03 3512.888889 1.90E-04 1.90E-04 1.90E-04 8_11:6.50E-03;8_5:2.25E-02;8_9:1.21E-02
+9 1.52E-02 2.34E-02 7.00E-03 3512.888889 2.63E-04 2.63E-04 2.63E-04 9_8:1.21E-02
+10 2.28E-02 3.38E-02 1.18E-02 3512.888889 4.45E-04 4.45E-04 4.45E-04 10_12:1.59E-02;10_7:8.76E-03
+11 4.44E-02 5.60E-02 3.28E-02 3512.888889 2.47E-04 2.47E-04 2.47E-04 11_14:5.62E-03;11_8:6.50E-03
+12 2.83E-02 1.90E-02 3.76E-02 3512.888889 1.62E-04 1.62E-04 1.62E-04 12_13:7.55E-03;12_10:1.59E-02;12_15:4.46E-03
+13 1.16E-02 1.46E-02 8.60E-03 3512.888889 1.70E-04 1.70E-04 1.70E-04 13_12:7.55E-03;13_14:6.55E-03
+14 4.60E-02 3.74E-02 5.46E-02 3512.888889 1.61E-04 1.61E-04 1.61E-04 14_11:5.62E-03;14_13:6.55E-03;14_16:5.07E-03
+15 2.01E-02 2.54E-02 1.48E-02 3512.888889 1.93E-04 1.93E-04 1.93E-04 15_18:5.09E-03;15_12:0.0044594326370586135
+16 1.28E-02 2.16E-02 4.00E-03 3512.888889 1.94E-04 1.94E-04 1.94E-04 16_19:0.01074071841628857;16_14:5.07E-03
+17 1.98E-02 3.18E-02 7.80E-03 3512.888889 3.48E-04 3.48E-04 3.48E-04 17_18:1.14E-02
+18 8.00E-02 9.62E-02 6.38E-02 3512.888889 1.93E-04 1.93E-04 1.93E-04 18_21:6.97E-03;18_15:5.09E-03;18_17:1.14E-02
+19 2.31E-01 2.33E-01 2.29E-01 3512.888889 2.49E-04 2.49E-04 2.49E-04 19_22:5.13E-03;19_20:8.25E-03;19_16:1.07E-02
+20 1.13E-01 3.26E-02 1.94E-01 3512.888889 1.81E-04 1.81E-04 1.81E-04 20_19:8.25E-03
+21 1.91E-02 2.80E-02 1.02E-02 3512.888889 3.70E-04 3.70E-04 3.70E-04 21_18:6.97E-03;21_23:7.72E-03
+22 2.33E-02 3.96E-02 7.00E-03 3512.888889 1.87E-04 1.87E-04 1.87E-04 22_19:5.13E-03;22_25:7.16E-03
+23 6.02E-02 7.44E-02 4.60E-02 3512.888889 2.62E-04 2.62E-04 2.62E-04 23_21:7.72E-03;23_24:9.50E-03
+24 2.17E-02 3.36E-02 9.80E-03 3512.888889 3.61E-04 3.61E-04 3.61E-04 24_23:9.50E-03;24_25:5.22E-03
+25 1.31E-02 1.92E-02 7.00E-03 3512.888889 1.42E-04 1.42E-04 1.42E-04 25_22:7.16E-03;25_26:4.25E-03;25_24:5.22E-03
+26 1.41E-02 1.90E-02 9.20E-03 3512.888889 1.65E-04 1.65E-04 1.65E-04 26_25:4.25E-03
+Table 2: Extended ibmq_mumber’s noise model including full detail of the gate error rate captured on 26-March-2024 from IBM.